Probability and Statistics I GSIA, Carnegie Mellon University

45-733, Spring 2002

Homework #5, due Thursday, February 21, 2002

Please show all relevant work! Please do the following problems from the book:

pp 263-5  1,11
pp 267-9  23
pp 289-92 1,15
pp 299-300 19,33

If you get stuck, it would be wise to try to work even-numbered problems similar to the ones assigned, since many of these have solutions in the back of the book.